A bowl contains five green olives and eight black ones. Will picks two olives. What is the probability that he gets one green and one black olive without replacement? Show work.
Answer by nerdybill(7384)

A bowl contains five green olives and eight black ones. Will picks two olives. What is the probability that he gets one green and one black olive without replacement?
first pick: 5/13
second pick: 8/12 = 2/3
.
Probability then is:
5/13 * 2/3 = 10/39 or 0.256 or 25.6%